# Sitcom-Crafter: A Plot-Driven Human Motion Generation System in 3D Scenes

### v1

@@ Abstract
Recent advancements in human motion synthesis have focused on specific types of motions, such as human-scene interaction, locomotion or human-human interaction, however, there is a lack of a unified system capable of generating a diverse combination of motion types. In response, we introduce Sitcom-Crafter, a comprehensive and extendable system for human motion generation in 3D space, which can be guided by extensive plot contexts to enhance workflow efficiency for anime and game designers. The system is comprised of eight modules, three of which are dedicated to motion generation, while the remaining five are augmentation modules that ensure consistent fusion of motion sequences and system functionality. Central to the generation modules is our novel 3D scene-aware human-human interaction module, which addresses collision issues by synthesizing implicit 3D Signed Distance Function (SDF) points around motion spaces, thereby minimizing human-scene collisions without additional data collection costs. Complementing this, our locomotion and human-scene interaction modules leverage existing methods to enrich the system’s motion generation capabilities. Augmentation modules encompass plot comprehension for command generation, motion synchronization for seamless integration of different motion types, hand pose retrieval to enhance motion realism, motion collision revision to prevent human collisions, and 3D retargeting to ensure visual fidelity. Experimental evaluations validate the system’s ability to generate high-quality, diverse, and physically realistic motions, underscoring its potential for advancing creative workflows.
@@ 3.1 System Overview
Before delving into our system, we recommend first reviewing Appendix A for an understanding of the methods closely related to our approach and some denotation symbols, which are placed there due to page length constraints. Fig. 2 demonstrates the entire workflow of Sitcom-Crafter. The initial input is a 3D scene with semantic information about its objects. The plot context can either be automatically generated based on the 3D scene information or manually designed by users. A language model is then employed to comprehend the plot context and transform it into recognizable commands for character motions. These commands are distributed to three motion generation modules responsible for crafting different types of motions: human locomotion, human-scene interaction, and human-human interaction. During the generation process, a motion synchronization module ensures consistency between synthesized frames from different generation modules. A hand pose retrieval module supplements hand poses to enhance the expressiveness of the human motions. To increase the realism and naturalness of the motions, a collision revision module adjusts the generated motions to prevent collisions between characters, and a motion retargeting module maps the generated motions to existing high-quality 3D human assets for improved visual fidelity.
@@ Appendix B More Details of Generation Modules
B.1 Human-Human Interaction Generation
@@
In contrast, marker points (Loper et al., 2014) can effectively represent body shape, and we can extract similarly positioned marker points from both SMPL and SMPL-X meshes. This approach enhances the network’s ability to consume training data from different sources. Our human-human interaction generator, therefore, uses marker points as body representation, both enabling the model to consume training data from different sources as well as aligning it with the representations used in the other two generation modules, GAMMA (Zhang & Tang, 2022) and DIMOS (Zhao et al., 2023). This results in a single unified body representation across the system.
@@
B.2 Human Locomotion and Human-Scene Interaction Generation
@@ C.2 Motion Synchronization
To ensure consistency between motions from different modules, we adopt the Slerp technique Shoemake (1985). This is primarily applied to transitions between other motions and human-human interactions, or vice versa, where there can be noticeable motion jumps despite conditioning the generation on the last frame (as mentioned in Section 3.2). Specifically, we designate the first and last four frames of generated human-human interactions as buffer areas for interpolation, using Slerp for rotation interpolation and linear interpolation for translation. This results in smoother motion transitions.
@@ Appendix H More Visual Results
Figures 22 and 23 present additional visual comparisons of different human-human interaction modules. Fig. 24 and Fig. 25 showcases further visual generation results guided by long plots.
@@ Appendix I Limitations and Outlooks
While the system excels in generating diverse human motion types, several constraints warrant future improvements. Firstly, the manual setup of camera poses for rendering sitcoms from generated human motion remains a bottleneck. Implementing an automatic camera-following network could streamline this process and enhance sitcom quality. Secondly, inherited limitations from DIMOS (Zhao et al., 2023) restrict human-scene interactions to sitting and lying scenarios. Expanding the repertoire to include a wider range of interactions requires access to high-quality, open-access human-scene interaction training data. Thirdly, due to differing training strategies and datasets for the three generative modules, there are still noticeable transitions between single human motions and human-human interactions. Additionally, in narrow spaces, physics constraints can prevent the generation of more dynamic interactions. Unifying the training setup and further enlarging human-human interaction training dataset may help address these challenges. Fourthly, the system currently focuses exclusively on human motions, neglecting movement of scene objects (Xu et al., 2023). Integrating basic, generalized object movements initially and gradually scaling to complex movements could address this gap. Lastly, the system is currently confined to motions within the same floor layer. Addressing interactions between different layers necessitates developing a dedicated generation module for seamless integration with the existing system.

### v2

@@ Abstract
Recent advancements in human motion synthesis have focused on specific types of motions, such as human-scene interaction, locomotion or human-human interaction, however, there is a lack of a unified system capable of generating a diverse combination of motion types. In response, we introduce Sitcom-Crafter, a comprehensive and extendable system for human motion generation in 3D space, which can be guided by extensive plot contexts to enhance workflow efficiency for anime and game designers. The system is comprised of eight modules, three of which are dedicated to motion generation, while the remaining five are augmentation modules that ensure consistent fusion of motion sequences and system functionality. Central to the generation modules is our novel 3D scene-aware human-human interaction module, which addresses collision issues by synthesizing implicit 3D Signed Distance Function (SDF) points around motion spaces, thereby minimizing human-scene collisions without additional data collection costs. Complementing this, our locomotion and human-scene interaction modules leverage existing methods to enrich the system’s motion generation capabilities. Augmentation modules encompass plot comprehension for command generation, motion synchronization for seamless integration of different motion types, hand pose retrieval to enhance motion realism, motion collision revision to prevent human collisions, and 3D retargeting to ensure visual fidelity. Experimental evaluations validate the system’s ability to generate high-quality, diverse, and physically realistic motions, underscoring its potential for advancing creative workflows. Project page: https://windvchen.github.io/Sitcom-Crafter.
[…]
@@ 3.1 System Overview
Before delving into our system, we recommend first reviewing Appendix A for an understanding of the methods closely related to our approach and some denotation symbols, which are placed there due to page length constraints. Fig. 2 demonstrates the entire workflow of Sitcom-Crafter. The initial input is a 3D scene with semantic information about its objects. The plot context can either be automatically generated based on the 3D scene information or manually designed by users. A language model is then employed to comprehend the plot context and transform it into recognizable commands for character motions. These commands are distributed to three motion generation modules for crafting different types of motions: human locomotion, human-scene interaction, and human-human interaction (for further details, refer to Appendix B.1). During the generation process, a motion synchronization module ensures consistency between synthesized frames from different generation modules. A hand pose retrieval module supplements hand poses to enhance the expressiveness of human motions. To increase the realism and naturalness of the motions, a collision revision module adjusts the generated motions to prevent collisions between characters, and a retargeting module maps the generated motions to existing high-quality 3D human assets for improved visual fidelity.
@@ Appendix B More Details of Generation Modules
Figure 7: Workflow within Generation Modules. Given the command lists derived from the Plot Comprehension module, each command is sequentially processed by the corresponding generation module based on its type and order. The motion generation of the current module is conditioned on the motion frames generated in the previous step, ensuring seamless chaining of motion segments. Note that the flowchart provides a simplified overview of the logic within the generation modules. For detailed explanations, please refer to the sections dedicated to each system module.
B.1 Detailed Workflow within Generation Modules
To clarify the overall logic of the generation modules used for synthesizing human motions and interactions, we illustrate the workflow using the example in Fig. 2, aiming to facilitate understanding. Consider the plot:“The scene opens with Jack, a young artist, sitting on the sofa. He stands up and paces nervously near the window. Amy, his girlfriend, enters the room, stretches out her arms and hugs him”. The system first employs the Plot Comprehension module to interpret the plot and generate two motion command lists. For Jack, the commands include: sit on sofa, stand up from sofa, walk near window, and finally, hug with Amy. For Amy, the commands include: position near the door, walk through door, and finally, hug with Jack. While these commands are illustrated in a conceptual and abstract manner for clarity, their actual format is detailed in Appendix B.3 and C.1).
Based on the type and order of the commands, the system invokes the appropriate generation modules. For Jack, it first calls the Human-Scene Interaction Generation Module (corresponding to stand up from sofa), then the Human Locomotion Generation Module (corresponding to walk near the window), and finally the Human-Human Interaction Generation Module (corresponding to hug with Amy). Similarly, for Amy, the system calls the Human Locomotion Generation Module (corresponding to walk through door) and then the Human-Human Interaction Generation Module (corresponding to hug with Jack). The initial poses of Jack (sitting on the sofa) and Amy (positioned near the door) are generated using either COINS (Zhao et al., 2022) or VPoser (Pavlakos et al., 2019) to establish the first frame of the human pose.
Regarding the chaining of motions across different modules, it is important to note that motions are not generated in parallel. Instead, they follow the order of commands in the motion command list: the first command is processed first, followed by subsequent commands in sequence. Each generation module conditions its outputs on frames generated by the previously invoked module. For example, the human locomotion and human-scene interaction modules use recurrent neural networks, enabling autoregressive generation of subsequent frames (refer to Appendix A), while the human-human interaction module explicitly conditions on the last frame (refer to Appendix B.2). This approach allows the system to generate long-duration motions aligned with the narrative guidance of extended plot contexts.
To better illustrate the workflow within the generation modules, a visual flowchart is provided in Fig. 7. It should be noted that we have simplified several technical details, such as handling frame length misalignment or addressing human collision issues during locomotion, to present a high-level overview of the logic within the generation modules. For a comprehensive explanation, please refer to the subsequent sections detailing each module of our system.
B.2 Human-Human Interaction Generation
@@
In contrast, marker points (Loper et al., 2014) can effectively represent body shape, and we can extract similarly positioned marker points from both SMPL and SMPL-X meshes. This approach enhances the network’s ability to consume training data from different sources. Our human-human interaction generator, therefore, uses marker points as body representation, both enabling the model to consume training data from different sources as well as aligning it with the representations used in the other two generation modules, GAMMA (Zhang & Tang, 2022) and DIMOS (Zhao et al., 2023). This results in a single unified body representation across the system. Beyond these benefits, such 3D-point based human representation, as validated by (Kolotouros et al., 2019; Zhang et al., 2021), captures richer body shape information and avoids periodicity and discontinuity issues in parametric rotation prediction, improving training and convergence efficiency.
@@
B.3 Human Locomotion and Human-Scene Interaction Generation
@@ C.2 Motion Synchronization
To ensure consistency between motions from different modules, we adopt the Slerp technique Shoemake (1985). This is primarily applied to transitions between other motions and human-human interactions, or vice versa, where there can be noticeable motion jumps despite conditioning the generation on the last frame (as mentioned in Section 3.2). Specifically, we designate the first and last four frames of generated human-human interactions as buffer areas for interpolation, using Slerp for rotation interpolation and linear interpolation for translation. This results in smoother motion transitions. It is worth noting that another potential solution for ensuring consistency is leveraging learning-based motion interpolation methods (Wang et al., 2021; Cohan et al., 2024), which can produce more natural human in-between motions. This approach could be seamlessly integrated into our system framework due to its modular design. However, we currently adopt the more efficient Slerp-based method and leave the exploration of learning-based solutions for future improvements.
@@ Appendix H More Visual Results
Figures 23 and 24 present additional visual comparisons of different human-human interaction modules. Fig. 25 and Fig. 26 showcases further visual generation results guided by long plots.
@@ Appendix I Limitations and Outlooks
While the system excels in generating diverse human motion types, several constraints warrant future improvements. Firstly, the manual setup of camera poses for rendering sitcoms from generated human motion remains a bottleneck. Implementing an automatic camera-following network could streamline this process and enhance sitcom quality. Secondly, inherited limitations from DIMOS (Zhao et al., 2023) restrict human-scene interactions to sitting and lying scenarios. Expanding the repertoire to include a wider range of interactions requires access to high-quality, open-access human-scene interaction training data. Thirdly, due to differing training strategies and datasets for the three generative modules, there are still noticeable transitions between single human motions and human-human interactions. Additionally, in narrow spaces, physics constraints can prevent the generation of more dynamic interactions. Unifying the training setup and further enlarging human-human interaction training dataset may help address these challenges. Fourthly, while the scene-aware human-human interaction module ensures physics alignment and prevents collisions, it does not significantly enrich the diversity or complexity of interaction motions themselves. The random synthesis of objects improves physical plausibility but does not fully capture the rich, entangled context of real-world human interactions. Generating more realistic and diverse human interaction motions will likely require exhaustive data collection efforts, such as motion capture or automated extraction from images and videos. Future advancements, possibly driven by scaling laws akin to those observed in models like Stable Diffusion (Rombach et al., 2022) or ChatGPT (OpenAI, 2022), could further benefit this area by leveraging massive datasets. Fifthly, the system currently focuses exclusively on human motions, neglecting movement of scene objects (Xu et al., 2023). Integrating basic, generalized object movements initially and gradually scaling to complex movements could address this gap. Lastly, the system is currently confined to motions within the same floor layer. Addressing interactions between different layers necessitates developing a dedicated generation module for seamless integration with the existing system.